# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Kwang-Ting Cheng1,2  Long Chen1†

### v2

@@ Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning. However, they remain limited in visual relation understanding, struggling even with binary relation detection, let alone N-ary relations involving multiple semantic roles. The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to unreliable outputs, hallucinations, and over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it). To this end, we propose Relation-R1, the first unified relation comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided supervised fine-tuning (SFT) and group relative policy optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-rewards optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Furthermore, we investigate the impact of various CoT strategies within this framework, demonstrating that a specific-to-general progressive approach in CoT guidance further improves generalization, especially in capturing synonymous N-ary relations. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
1 Introduction
@@
Recent advancements in MLLMs have significantly enhanced holistic image understanding [1, 2] and object-level grounding [3, 4, 5, 6, 7] capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding — a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
@@
Figure 1: Illustrations of our Relation-R1 and other MLLMs. (a) Grounded MLLMs: Link objects in text to image regions but fail to capture object relations. (b) Relation Comprehension MLLMs: Model pairwise object relations but lack generalization to N-ary interactions. (c) Unified Relation Comprehension MLLMs: Jointly handle binary and $N$-ary relation detection.
@@
Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models [8, 9] struggle to capture N-nary relationships. Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) [10, 11]. The neglect of these crucial structural semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, may result in suboptimal relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to both hallucinations and over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions). Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
@@
Fortunately, DeepSeek-R1-Zero [12] has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. However, due to the inherent difficulty of reference models in understanding structural relationships, the direct application of RL often struggles to produce consistently formatted outputs that are claimed in the input prompt [13]. Conversely, when solely adopting supervised fine-tuning, the model suffers from poor generalization due to its overfitting to fixed training patterns and the limited diversity of annotated relational data.
To this end, inspired by the breakthrough of R1-series [12, 14, 15] in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
@@
Stage 2. RL: Refines the structured outputs generated by the SFT model through group relative policy optimization. The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process [12]; 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity. This stage employs policy gradient updates to cultivate the model to explore various potential solutions and optimizes its output based on defined multi-rewards, fostering robust relational reasoning and generalization.
@@
We validate the effectiveness of Relation-R1 on both the binary relation comprehension task (e.g., scene graph generation [16, 8]) and N-ary relation comprehension task (e.g., grounded situation recognition [10]) across diverse experimental settings. Experiments on the PSG [17] dataset and SWiG [10] dataset demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are fourfold:
@@
  • Experimental results on the prevalent relation understanding datasets (e.g., +6.84$\sim$6.90gains on PSG dataset) under different settings demonstrate the effectiveness of Relation-R1.
2 Related Work
@@
Reasoning in MLLMs. Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks [18, 12]. A pivotal breakthrough, DeepSeek-R1 [12], leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards [14, 19, 20] For multimodal mathematical reasoning, frameworks such as R1-Onevision [14], Virgo [19], and MM-Eureka [21] jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 [20] combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT [22] and R1-Omni [15] utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero [23] employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, Relation-R1 explicitly integrates relational reasoning into MLLMs via RL, enabling systematic inference for both scene interpretation and event prediction.
@@
Scene Graph Generation (SGG). SGG is a critical task in scene understanding, with prior approaches falling into two categories: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation [24, 25, 26, 27]; 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods [28, 29, 30]) but often lacks multi-role involved relation modeling. Recent efforts in open-vocabulary SGG (OVSGG) leverage Vision-Language Models (VLMs) [31, 32, 33] or MLLMs [8, 9, 34] to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT. Our Relation-R1 addresses OVSGG challenges by integrating a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
@@
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al. [10] pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works [11, 35, 36, 37] improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution [24, 25, 26], potentially limiting real-world applicability. While LEX [38] realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R1, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
3 Methodology
@@
Figure 2: Overview of the proposed Relation-R1 framework. (a) SFT with progressive guidance of cognitive CoT, which enforces task-specific reasoning processes and format alignment. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-nary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
3.1 Preliminaries
@@
Binary Relation Detection. This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects [8]. The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates in < box> tags to localize the entity in the image, e.g., <ref>person</ref><box>[[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes.
N-ary Relation Detection. Similarly, this task aims to output a open-ended grounded situation frame [10] structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates. Each role explicitly defines the entity’s participation in the activity, e.g., <agent>person</agent><box>[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]</box>.
3.2 Supervised Fine-Tuning
@@
As discussed in §1, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs. To address this, we employ SFT as a foundational stage to establish task-specific output format adherence (c.f., Figure 2(a)).
@@
However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures [13]. To mitigate this potential issue, we introduce the cognitive chain-of-thought, a methodology that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference. These cognitive CoTs are explicitly encapsulated within <think> tags, during SFT. This ensures that the model retains its capacity for multi-step reasoning while learning to produce structured outputs conforming to the desired format.
@@
MLLM-generated CoT. For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL [39]) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth scene graph of the current image, and a CoT generation instruction. The former one makes the MLLM generate the correct task format. Meanwhile, the latter two anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed cognitive CoT generation prompt is left in the appendix §B.
Progressive CoT Guidance. With a progressive paradigm that involves initial training on easy-to-learn fully template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains the foundational capability to generate format-compliant outputs and grounded reasoning processes, which is critical for GRPO training.
3.3 Group Relative Policy Optimization
Following the Deepseek-R1 [12], the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods [40], GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity [12]. Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{1},o_{2},\ldots,o_{G}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging the following objectives:
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\epsilon,1+\epsilon)A_{i}\right)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\right],$ (1)
@@ 3.3 Group Relative Policy Optimization
$\small A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy is formulated as:
$\small D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
@@
The reward $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-nary relation rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
@@
$\small r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
Binary Relation Rewards. Inspired by established evaluation practices [17, 8], we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample. The former is the ratio of correctly predicted triplets to the total ground-truth triplets in a sample; the latter is the average recall across all predicate categories. Specifically, a triplet is deemed correct if:
@@
$\small r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
@@
N-ary Relation Rewards. Similarly, to evaluate higher-order relational structures in grounded situation frames, we introduce N-ary relation rewards, which assess the model’s ability to detect primary activity (verb) and multi-entity interactions. Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample [10]. These respectively represent the correct proportion in predicted entity categories and roles, and the spatial localization of those entities:
  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the ground truth, and the predicted activity matches the ground-truth.
  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the ground-truth’s bounding box.
@@
$\small r_{\text{n-ary}}(o)=\beta\cdot V+(1-\beta)\cdot V_{grnd},$ (6)
@@
4 Experiment
4.1 Experimental Setup
Datasets. 1) Panoptic Scene Graph (PSG): The official PSG dataset [17], built upon the COCO dataset [41], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing. It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories. Notably, ASMv2 [8] reconstructs the dataset into a scene graph caption format, utilizing a split of 42,250 training images and 1,000 evaluation images. To mark fair comparison, we followed ASMv2 [8], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training. Besides, we also present results using the conventional scene graph format, i.e., a list of visual triplets structured as [subject, subject bounding box, object, object bounding box, relation] on the official dataset [^1] More details of the scene graph format are in the appendix §C.. 2) SWiG: The SWiG dataset [10] extends the imSitu dataset [42] by adding extra bounding box annotations for semantic roles, with 69.3% of semantic roles localized. Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators. The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories. Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
Evaluation Metrics. For binary relation detection, we followed the standard SGG evaluation protocol [24, 8] 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate categories. As for N-ary relation detection, we adopted the same five evaluation metrics as [10]: 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
Implementation Details. Due to space constraints, details are provided in the appendix §A.
@@ 4.1 Experimental Setup
         Method Model Size Format Recall mRecall Mean
@@
IMP [16] ${}_{\text{CVPR'17}}$ 16.50 6.50 11.50
MOTIFS [25] ${}_{\text{CVPR'18}}$ 20.00 9.10 14.55
VCTree [43] ${}_{\text{CVPR'19}}$ 20.60 9.70 15.15
GPSNet [44] ${}_{\text{CVPR'20}}$ 17.80 7.00 12.4
PSGFormer [17] ${}_{\text{ECCV'22}}$ 18.60 16.70 17.65
@@
TextPSG [45] ${}_{\text{ICCV'23}}$ Standard Scene Graph 4.80 - -
R1-SGG [46] ${}_{\text{ArXiv'25}}$ 2B 27.83 17.03 22.43
R1-SGG [46] ${}_{\text{ArXiv'25}}$ 7B 28.77 17.55 23.16
@@
ASMv2 [8] ${}_{\text{ECCV'24}}$ 13B Scene Graph Caption 14.20 10.30 12.23
LLaVA-SpaceSGG [9] ${}_{\text{WACV'25}}$ 13B 15.43 13.23 14.33

### v3

# Relation-R1: Progressively Cognitive Chain-of-Thought Guided Reinforcement Learning for Unified Relation Comprehension

@@ Abstract
Recent advances in multi-modal large language models (MLLMs) have significantly improved object-level grounding and region captioning. However, they remain limited in visual relation understanding, struggling even with binary relation detection, let alone N-ary relations involving multiple semantic roles. The core reason is the lack of modeling for structural semantic dependencies among multi-entities, leading to over-reliance on language priors (e.g., defaulting to “person drinks a milk” if a person is merely holding it). To this end, we propose Relation-R1, the first unified relation comprehension framework that explicitly integrates cognitive chain-of-thought (CoT)-guided supervised fine-tuning (SFT) and group relative policy optimization (GRPO) within a reinforcement learning (RL) paradigm. Specifically, we first establish foundational reasoning capabilities via SFT, enforcing structured outputs with thinking processes. Then, GRPO is utilized to refine these outputs via multi-rewards optimization, prioritizing visual-semantic grounding over language-induced biases, thereby improving generalization capability. Furthermore, we investigate the impact of various CoT strategies within this framework, demonstrating that a specific-to-general progressive approach in CoT guidance further improves generalization, especially in capturing synonymous N-ary relations. Extensive experiments on widely-used PSG and SWiG datasets demonstrate that Relation-R1 achieves state-of-the-art performance in both binary and N-ary relation understanding.
Codegithub.com/HKUST-LongGroup/Relation-R1
Introduction
@@
Recent advancements in MLLMs have significantly enhanced holistic image understanding (liu2023visual; zhu2024minigpt) and object-level grounding (youferret; zhang2024llava; rasheed2024glamm; yuan2024sr; yuan2025voxel; lai2024lisa; peng2024kosmos) capabilities, enabling tasks like region captioning and referring question answering. However, current MLLMs exhibit critical limitations in relational scene understanding – a core competency for achieving human-like visual cognition, as illustrated in Figure 1(a).
@@
Figure 1: Illustrations of our Relation-R1 and other MLLMs. (a) Grounded MLLMs: Link objects in text to image regions but fail to capture object relations. (b) Relation Comprehension MLLMs: Model pairwise object relations but lack generalization to N-ary interactions. (c) Unified Relation Comprehension MLLMs: Jointly handle binary and NN-ary relation detection.
@@
Despite notable progress in binary relation detection, i.e., identifying subject-object interactions (e.g., “child-drinking-glass”), current models (wang2024all; llava_spacesgg2025) struggle to capture N-ary relationships. Accurately modeling such relationships requires models to identify multiple N-th entities engaged in the relation with their distinct semantic roles (what, who, and where concepts) (pratt2020grounded; cho2021grounded). The neglect of these crucial structural semantic dependencies among multi-entities, such as the glass’s role as the functional container of milk in the drinking action, may result in suboptimal relation triplets, e.g., “child-drinking-glass”. Furthermore, these unreliable and superficial pair-wise relations may make the MLLMs prone to over-reliance on language priors rather than visual-semantic cues (e.g., evidence from semantically grounded visual regions). Consequently, there arises an urgent demand for a unified framework capable of addressing these shortcomings. This raises a critical question: How can we devise a unified framework for joint binary and N-ary relational reasoning with grounded cues?
@@
Fortunately, DeepSeek-R1-Zero (DeepseekR1) has successfully demonstrated the emergence of reasoning capabilities in LLMs purely through RL. However, due to the inherent difficulty of reference models in understanding structural relationships, directly applying RL often struggles to produce consistently formatted outputs that are claimed in the input prompt (chu2025sft). Conversely, when solely adopting supervised fine-tuning, the model suffers from poor generalization due to its overfitting to fixed training patterns and the limited diversity of annotated relational data.
To this end, inspired by the breakthrough of R1-series (DeepseekR1; yang2025r1onevision; zhao2025r1) in enabling reasoning and the ability of SFT to provide structured guidance, we propose a unified two-stage relation comprehension framework, Relation-R1, which combines the advantages of both supervised fine-tuning and reinforcement learning to empower MLLMs with relational reasoning and generalization capability:
@@
Stage 2. RL: Refines the structured outputs generated by the SFT model through group relative policy optimization. The learning process is guided by multi-rewards: 1) format rewards, which incentivize the generation of thinking process (DeepseekR1); 2) binary relation rewards, which encourage accurate identification of pairwise relationships between entities; and 3) N-ary relation rewards, which promote the comprehensive understanding of more complex, multi-entity activity. This stage employs policy gradient updates to cultivate the model to explore various potential solutions and optimizes its output based on defined multi-rewards, fostering robust relational reasoning and generalization.
@@
We validate the effectiveness of Relation-R1 on both binary relation comprehension (e.g., scene graph generation (xu2017scene; wang2024all)) and N-ary relation comprehension task (e.g., grounded situation recognition (pratt2020grounded)) across diverse experimental settings. Experiments on the PSG (yang2022panoptic) and SWiG (pratt2020grounded) datasets demonstrate state-of-the-art performance in relational reasoning while maintaining generalization in ambiguous contexts. In summary, our contributions are fourfold:
@@
[…]
Related Work
@@
Reasoning in MLLMs. Recent LLMs have demonstrated advanced reasoning capabilities by emulating human-like stepwise thinking, significantly enhancing performance on complex tasks (jaech2024openai; DeepseekR1). A pivotal breakthrough, DeepSeek-R1 (DeepseekR1), leverages large-scale RL with formatting and result-oriented rewards to enable LLMs to autonomously generate human-like complex Chain-of-Thought (CoT) reasoning, achieving state-of-the-art results across diverse domains. MLLMs extend these capabilities through systematic integration of cross-modal knowledge representation and task-specific rewards (yang2025r1onevision; du2025virgo; huang2025vision; shen2025tarpro) For multimodal mathematical reasoning, frameworks such as R1-Onevision (yang2025r1onevision), Virgo (du2025virgo), and MM-Eureka (meng2025mm) jointly reason over visual and textual inputs to solve quantitative tasks, while Vision-R1 (huang2025vision) combines cold-start initialization with GRPO to refine complex CoT reasoning. For fine-grained classification and grounding tasks, Visual-RFT (liu2025visual) and R1-Omni (zhao2025r1) utilize GRPO-based reinforcement algorithms and verifiable rewards to enhance contextual reasoning. As for the pixel-level understanding task, SegZero (liu2025seg) employs reinforcement learning to achieve high-resolution semantic segmentation. Despite these advances, existing MLLMs predominantly focus on object-centric recognition rather than unified visual relation understanding, which requires compositional reasoning about geometric, semantic, and functional interactions between objects and contexts. To bridge this gap, Relation-R1 explicitly integrates relational reasoning into MLLMs via RL, enabling systematic inference for both scene interpretation and event prediction.
@@
Scene Graph Generation (SGG). SGG is a critical task in scene understanding, with prior approaches falling into two groups: 1) Two-stage SGG, which sequentially detects objects and infers pairwise relations but suffers from error propagation (tang2020unbiased; zellers2018neural; li2022devil; li2023compositional; shi2025easy); 2) One-stage SGG, which unifies detection and relation prediction in end-to-end frameworks (e.g., DETR-based methods (carion2020end; li2022sgtr; chen2020simple)) but lacks multi-role involved relation modeling. Recent efforts in open-vocabulary SGG (OVSGG) employ Vision-Language Models (VLMs) (chen2024scene; li2023zero; li2025inter; chen2024expanding) or MLLMs (wang2024all; llava_spacesgg2025; li2024pixels) to handle novel entities/relations, yet limited modeling of complex multi-entity interactions, and suffer from overfitting in base categories during SFT. Our Relation-R1 addresses OVSGG challenges by adopting a cognitive CoT-guided RL framework to model complex multi-entity interactions and mitigate SFT overfitting, enabling robust zero-shot reasoning without predefined category constraints.
@@
Grounded Situation Recognition (GSR). GSR extends scene understanding by jointly modeling actions (verbs) and their associated semantic roles. Pratt et al(pratt2020grounded) pioneered this field with a two-stage RNN-based framework: verb detection followed by noun localization. Subsequent works (cho2021grounded; wei2021rethinking; cho2022collaborative; cheng2022gsrformer) improved this pipeline by integrating transformers and semantic relation modeling, enabling more coherent scene interpretations. Despite these efforts, previous efforts that depend on annotated training samples may face challenges like long-tailed distribution (tang2020unbiased; li2024nicest; shao2024knowledge; chen2023addressing), potentially limiting real-world applicability. While LEX (lei2024seeing) realizes training-free zero-shot GSR, it fails to provide an end-to-end framework and exhibits suboptimal performance and computational efficiency due to its reliance on LLMs’ generated descriptions. This paper proposes Relation-R1, a unified framework that bridges the gap between open-ended capabilities and end-to-end training.
@@
Figure 2: Relation-R1 framework. (a) SFT with progressive guidance of cognitive CoT, which enforces task-specific reasoning processes and format alignment. (b) RL with GRPO, employs rule-based rewards (e.g., format rewards, binary relation rewards, and N-ary relation rewards) to cultivate robust relation reasoning and generalization ability through policy gradient updates.
Methodology
@@
Preliminaries. 1) Binary Relation Detection: This task requires the model to generate an open-ended scene graph caption given an image, explicitly marking semantic relationships between pairwise objects (wang2024all). The output is a triplet-contained caption annotated with formal tags: objects are denoted by <ref> tags followed by their bounding box coordinates (top-left/bottom-right) in < box>, e.g., <ref>person</ref><box>$[$$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]]</box>. Predicates are enclosed in <pred> tags, followed by references to their subject and object entities via their corresponding bounding boxes. 2) N-ary Relation Detection: It aims to output a open-ended grounded situation frame (pratt2020grounded) structured as a primary activity (i.e., multi-entity involved relation), followed by a sequence of <role> tags along with corresponding bounding box coordinates. Each role explicitly defines the entity’s participation in the activity, e.g., <agent>person</agent><box>[$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]</box>.
Supervised Fine-Tuning
@@
As discussed above, directly applying GRPO encounters challenges in generating consistently formatted scene graph captions and grounded situation frame outputs. To address this, we employ SFT as a foundational stage to establish task-specific output format adherence (c.f., Figure 2(a)).
@@
However, relying solely on the target answer format during SFT risks compromising the model’s underlying reasoning capabilities by encouraging a superficial alignment with formulated structures (chu2025sft). To mitigate this potential issue, we introduce the cognitive chain-of-thought that integrates stepwise cognitive processes such as object classification, object grounding, and visual relationship inference. These cognitive CoTs are explicitly encapsulated within <think> tags, during SFT. This ensures that the model retains its capacity for multi-step reasoning while learning to produce structured outputs conforming to the desired format.
@@
MLLM-generated CoT. For the general reasoning process, we design a cognitive CoT generation prompt that explicitly makes an extra strong MLLM (e.g., Qwen 2.5-VL (Qwen2.5-VL)) to generate step-by-step reasoning processes. To be specific, this cognitive CoT generation prompt consists of a task definition, a ground-truth (GT) scene graph of the current image, and a CoT generation instruction. The former one makes the MLLM generate the correct task format. Meanwhile, the latter two anchors reasoning content to visual evidence, ensuring that the MLLM’s cognitive CoT aligns with both semantic validity and visual grounding rather than relying on category priors. The detailed cognitive CoT generation prompt is left in the Appendix.
Progressive CoT Guidance. With a progressive paradigm that involves initial SFT training on easy-to-learn template-based CoTs and subsequent fine-tuning on a small number of MLLM-generated CoTs, the MLLM gains foundational capability to generate format-compliant outputs and grounded reasoning processes, contributing to GRPO training.
Group Relative Policy Optimization
Following the Deepseek-R1 (DeepseekR1), the group relative policy optimization is adopted as the RL algorithm for optimizing our Relation-R1 framework, as illustrated in Figure 2(b). Distinct from critic-dependent methods (schulman2017proximal), GRPO directly calculates advantage estimates through response group comparisons sampled from the policy model, inherently eliminating the need for an extra critic network and reducing computational complexity (DeepseekR1). Key components include the a question $q$, policy model $\pi_{\theta}$, response group $\{o_{1},o_{2},\ldots,o_{G}\}$, and a frozen reference model $\pi_{\theta_{\text{ref}}}$, which acts as a regularization baseline to preserve prior knowledge. The GRPO objective function $J_{\text{GRPO}}(\theta)$ balances reward maximization and policy stability by leveraging the following objectives:
$J_{\text{GRPO}}(\theta)=\mathbb{E}_{q\sim\mathcal{Q},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}}$ (1)
$[\frac{1}{G}\!\sum_{i=1}^{G}\!\min\left(\!\rho_{i}A_{i},\text{clip}(\rho_{i},1\!-\!\epsilon,1\!+\!\epsilon)A_{i}\!\right)\!-\!\beta\!D_{\text{KL}}(\!\pi_{\theta}\!\|\!\pi_{\text{ref}})],$
@@
$A_{i}=\frac{r_{i}-\text{mean}\left(\{r_{1},\ldots,r_{G}\}\right)}{\text{std}\left(\{r_{1},\ldots,r_{G}\}\right)},$ (2)
here $r_{i}$ denotes the reward for response $o_{i}$. The KL divergence term enforces proximity to the reference policy:
$D_{KL}(\pi_{\theta}\|\pi_{\text{ref}})=\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}-\log\left(\frac{\pi_{\text{ref}}(o_{i}|q)}{\pi_{\theta}(o_{i}|q)}\right)-1.$ (3)
@@
The reward $r_{i}$ is formulated to align with the model’s learning objective for the unified relation understanding. It incorporates rule-based rewards, which consist of three crucial components: Format compliance rewards, binary relation rewards, and N-ary relation rewards. The former enforces syntactic validity by penalizing outputs that deviate from predefined structural constraints. The latter two components focus on task-specific performance.
@@
$r_{\text{form}}(o_{i})=\begin{cases}1&\text{if }o_{i}\text{ adheres to the format},\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
Binary Relation Rewards. Inspired by established evaluation practices (yang2022panoptic; wang2024all), we devise the sample-level triplet recall $R$ and sample-level mean recall $mR$ in the response of each sample. The former is the ratio of correctly predicted triplets to the total ground-truth triplets in a sample; the latter is the average recall across all predicate categories. Specifically, a triplet is deemed correct if:
@@
$r_{\text{binary}}(o)=\alpha\cdot R+(1-\alpha)\cdot mR,$ (5)
@@
N-ary Relation Rewards. Similarly, to test higher-order relational structures in grounded situation frames, we devise N-ary relation rewards to assess the model’s ability to detect primary activity (verb) and multi-entity interactions. Specifically, we calculate the entity value $V_{e}$ and grounded value $V_{grnd}$ for each sample (pratt2020grounded). They represent the correct proportion in predicted entity classes and roles, and the spatial localization of those entities, respectively:
  • Entity Accuracy: An entity within an event is considered correct if its predicted category and semantic role align with the GT, and the predicted activity matches the GT.
  • Spatial Localization: The predicted bounding box of each role obtains IoU >=0.5 with the GT’s bounding box.
@@
$r_{\text{n-ary}}(o)=\beta\cdot V_{e}+(1-\beta)\cdot V_{grnd},$ (6)
@@
Experiment
@@
  Method Size Recall mRecall Mean
@@
IMP (xu2017scene) - 16.50 6.50 11.50
MOTIFS (zellers2018neural) - 20.00 9.10 14.55
VCTree (tang2019learning) - 20.60 9.70 15.15
GPSNet (lin2020gps) - 17.80 7.00 12.4
PSGFormer (yang2022panoptic) - 18.60 16.70 17.65
@@
TextPSG (Zhao_2023_ICCV) - 4.80 - -
R1-SGG (chen2025compile) 2B 27.83 17.03 22.43
R1-SGG (chen2025compile) 7B 28.77 17.55 23.16
@@
ASMv2 (wang2024all) 13B 14.20 10.30 12.23
SpaceSGG (llava_spacesgg2025) 13B 15.43 13.23 14.33